OpenStudy (anonymous):

prove; sinxcosx(tanx+cotx)=1

OpenStudy (anonymous):

\[ab(\frac{b}{a}+\frac{a}{b})\]

OpenStudy (anonymous):

you get \[b^2+a^2\] or in this case \[\sin^2(x)+\cos^2(x)=1\]